# References.bib

### v1

@@ 4 Existing Literature on BC4LLM
Independently, the fields of both LLMs and blockchain research have grown substantially over the past several years. It is no surprise that the literature surrounding these topics has begun to morph and relate to each other. In previous research, we have seen LLMs for Blockchain Security [he_large_2024] as well as an introduction to the term BC4LLM in Luo et al [luo_bc4llm_2023] where they provide a comprehensive survey of blockchain for LLMs. However, they do not acknowledge the multitude of safety and security solutions that blockchain provides for certain LLM vulnerabilities. Effectively, Luo et al. [luo_bc4llm_2023] aims to introduce BC4LLM for trusted AI, enabling reliable learning corpora, secure training processes, and identifiable generated content. In juxtaposition, our survey aims to analyze possible BC4LLM solutions closely related to our definitions of safety (2) and security (1) when looking at inherent system vulnerabilities in LLMs. To begin our analysis, we define these security problems based on previous work and highlight areas of research that are applicable to areas of BC4LLM safety and security.
4.1 Blockchain for Large Language Model Security
@@
Few papers and experiments analyze how the integration of these two technological powerhouses interact with one another. We have seen benefits of this integration that apply to our definition of security (1). Balija et al. [balija_building_2024] introduced a peer-to-peer (P2P) federated LLM, namely PageRank, which works with a blockchain. This system operates in a fully decentralized capacity. Demonstrably, the blockchain implementation led to more efficient accuracy and latency results. With that being stated, Balija et al. [balija_building_2024] provide a developing direction in the field of BC4LLM to enhance system security. Below, we address several current vulnerabilities in LLMs and analyze them individually. In order to better understand these security problems, we categorize these vulnerabilities to their respective LLM training stages, highlight blockchain for AI works, and provide well-researched blockchain applications as a solution.
@@
Vulnerabilities are present at each step in the process of developing a LLM. In early methods of model training, we encounter adversarial attacks such as data poisoning and backdoor attacks within the corpus [shayegani_survey_2023, yan_protecting_2024]. Progressing into model fine-tuning and general use, the LLM can fall victim to prompt-based attacks, [abdali_securing_2024][zhao_prompt_2023] [yan_protecting_2024][liu_prompt_2024], inference attacks [hu_membership_2022] [keshk_privacy-preserving-framework-based_2020], and RAG related attacks [xue_badrag_2024] [cheng_trojanrag_2024][anderson_is_2024][deng_pandora_2024][zou_poisonedrag_2024] [zeng_good_2024]. These attacks are common vulnerabilities in both LLMs and AI since LLMs and AI are closely related as seen in Figure 2. Some of the threats against LLMs can be addressed by implementations from blockchain for AI (BC4AI) research, as elaborated below in Section 4.1.3. Considering the volume of potential attacks against LLMs, we make a further distinction of solutions that are specifically related to BC4LLM research and other blockchain-based solutions from BC4AI research. With this, we are able to highlight shared vulnerabilities for LLMs and AI. We provide an analysis of how blockchain can help defend against and mitigate these vulnerabilities, starting with threats during each phase of LLM training and utilization, continuing onto different blockchain solutions for AI inherent threats, and lastly noteworthy technology inherent attacks such as denial of service (DDoS) attacks and issues with supply chain logistics.
4.1.1 Blockchain for Threats within LLM Training Process
@@
Situated at a critical juncture in the process of model development is the issue of data selection. An area of paramount concern for this data is exactly how we can verify that training data is authentic and safe as well as tamper-proof from data poisoning attacks. A new approach may be the LLM’s ability to unlearn this poisoned data, or data that is unsafe according to our definition of safety 2. From Zuo et al. [zuo_federatedtrustchain_2024] they establish federated TrustChain as a method of enhancing LLM training and unlearning through a blockchain-based federated learning framework. By integrating blockchain using Hyperledger, their findings present that the framework is efficient in unlearning capabilities, showcasing an accuracy score initially of 99.15% and after unlearning, the accuracy drops to 0.70% [zuo_federatedtrustchain_2024]. This highlights the potential for BC4LLM to improve security and privacy guarantees, where the LLM can selectively forget specified data points while simultaneously preserving their performance with the leverage of Low-Rank Adaptation (LoRA) and tuning hyper-parameters. This method of a blockchain-enabled federated unlearning process is further detailed as a future research possibility that has been thoroughly explored by few, as emphasized later in Section 7.1.
@@
Another significant problem in the area of adversarial threats against LLMs is data poisoning. Gong et al. [gong_dynamic_2023] proposed a possible blockchain solution, introducing dynamic large language models (DLLM) on blockchains. Instead of using the traditional centralized data sets that LLMs are provided with, developing LLMs on blockchains enables the creation of decentralized datasets. These datasets are less likely to be tampered with and can be easily audited for accuracy. Gong et al. presents DLLM to evolve after the training process. This was implemented through neural network parameters that offer the LLM the ability to learn during its usage.
@@
Figure 4: We provide an overview of a two-level framework consisting of: a privacy-based blockchain that uses the secure hash algorithm 512 (SHA512) to generate hashes for data integrity. Then, the enhanced proof of work (ePoW) is used to authenticate data records and prevent data poisoning attacks from altering original data. These hashes of data blocks are linked to each other, called a Hash Chain. Then, for the second level, privacy-based variational autoencoder (VAE) for secure data transformation, ensures robust protection against inference attacks while maintaining the utility model for anomaly detection.
@@
Additionally, blockchain-based systems can help assess where data poisoning may occur and as shown in [keshk_privacy-preserving-framework-based_2020] blockchain can protect datasets and detect potential inference attacks through a two-level privacy preserving module. This research included a framework based on blockchain and deep learning, including two levels of privacy mechanisms as shown in Figure 4. For the first level, Keshk et al. [keshk_privacy-preserving-framework-based_2020] used SHA512 to generate secure hashes and then implemented an enhanced-proof-of-work (ePoW) technique for authenticating data records and preventing data poisoning attacks. The second level consisted of a VAE technique for converting original data into an encoded format for mitigating inference attacks that could be learned from system-based machine learning. In their testing, these mechanisms were effective in preventing data poisoning and inference attacks from manipulated smart power network datasets. BC4LLMs could benefit from this similar type of implementation, working with secure methods of hashing and blockchain-based deep learning privacy preservation techniques. By integrating a two-level privacy-preserving module, BC4LLMs can ensure data integrity and confidentiality while effectively detecting and mitigating both data poisoning and inference attacks.
@@
Poisoned data has been an interest with RAG in particular. For example, Xue et al. [xue_badrag_2024] developed a way to identify security vulnerabilities from a poisoned corpus, but they do not use blockchain as a solution. We address the absence of research on blockchain and RAG, especially when using blockchain to help prevent RAG security issues. We discuss this as a possible future research direction as there remains a current gap in research of blockchain-based RAG systems and elaborate on this topic in 7.2. Poisoned data overall is a major concern within LLMs and we offer blockchain as a potential source of ground truth to aid in mitigating this threat during the pre-training stages of LLMs and potentially mitigate RAG security concerns. In addition to data poisoning, LLMs are susceptible to backdoor attacks hidden in the training data during the LLM pre-training phase. Notable research by Zhao et al. [zhao_prompt_2023] introduced ProAttack which improves the stealth of backdoor attacks by accurately labeling poisoned data samples. As these attacks improve and become more sophisticated, it is crucial to explore robust defense mechanisms for LLMs. Few defense mechanisms using blockchain have been made, but we see in Li et al. [li_blockchain-based_2023] a proposal for a blockchain-based federated-learning framework (DBFL) that withstands backdoor attacks in a blockchain environment through incorporating an RLR aggregation strategy into the aggregation algorithm of a user and the addition of gradient noise to limit the effectiveness of backdoor attacks. The robustness of FL against backdoor attacks is enhanced by using various blockchain functions, including digital signature verification and simulation of chain resynchronization [li_blockchain-based_2023].
4.1.2 Blockchain for Threats within LLM Prompting and Utilization
@@
LLMs undergo continuous training through instruction tuning, alignment tuning, and fine-tuning, which warrants a list of vulnerabilities such as prompt injection [mbula_mboma_assessing_2023][shen_large_2023] and backdoor attacks [li_blockchain-based_2023]. These adversarial attacks are included under the term active privacy 3 where a malicious user attempts to gain unauthorized access. Blockchain technology, with its inherent transparency and immutability, holds the potential to mitigate and defend against these vulnerabilities. For instance, blockchain’s ability to ensure data integrity and provide traceability can play a pivotal role in defending against prompt injection attacks. Mbula et al. [mbula_mboma_assessing_2023] produce an overview of LLMs for blockchain and note the capabilities of blockchain, especially the transparency and immutability to provide a reliable audit trail of transactions to track and investigate any suspicious activities. Although not specifically focused on prompt injection, this approach showcases how blockchain can supplement security by offering a clear and immutable record of interactions. Applying this to BC4LLMs can help prevent suspicious users from continuously interacting with a LLM, allowing for traceability to stop the user from entering malicious prompts. We recognize prompt injection is a critical vulnerability in LLMs and AI-related systems, yet as noted in the survey from Shen et al. [shen_large_2023] few blockchain defenses for prompt injection are present in the current field of research.
@@
Inference attacks are also of critical importance to LLMs and to active privacy, where a malicious user tries to extract data from the LLM. The malicious user could also try to gain information about the training data, hence why the Taxonomy 3 has inference attacks standalone. As discussed previously in Section 4.1.1 from Keshk et al. [keshk_privacy-preserving-framework-based_2020] that uses Blockchain and DL to privacy preserve and prevent inference attacks through a framework as depicted in Figure 4. For more inference attack applicable work, a survey [hu_membership_2022] thoroughly discusses membership inference attacks on ML and provides a group of defenses including differential privacy, regularization, confidence masking, and knowledge distillation. In other works, there are instances of blockchain-based differential privacy methods [zhao_blockchain-based_2021][han_blockchain-based_2020][park_blockchain-based_2021], but current research that uses blockchain-based differential privacy frameworks to prevent inference attacks is limited; It is worth noting that the theoretical foundation and potential synergies of this combination are promising. Another area with limited research is blockchain as a defense for jailbreaking attacks, which exploit the inherent capabilities of LLMs to bypass restrictions. There are multiple articles defending LLMs from jailbreaking attacks, yet little to none fully include blockchain to prevent jailbreaking. Hu et al. [hu_blockchain-based_2021] explores a blockchain defense mechanism for malware checking on operating systems, indicating a possible direction for future research in integrating blockchain to defend against jailbreaking in LLMs. As previously explained in Section 3.2, backdoor attacks after the model has been trained are based on modifications to the trained model. The key blockchain-based federated-learning framework from Li et al. [li_blockchain-based_2023] discussed in detail in Section 4.1.1 used a combination of a blockchain environment and an RLR aggregation strategy to defend against backdoor attacks. This framework effectively coordinated FL processes and maintained learning security and user privacy. When testing backdoor attacks caused by malicious participants, the accuracy of the model increased when using the RLR aggregation strategy [li_blockchain-based_2023]. Given these findings, the possibility of leveraging blockchains transparency and immutability presents a robust mechanism for improving LLM security against active privacy threats. However, comprehensive integration and empirical validation of blockchain-based defenses in LLMs remain imperative to advance the field of BC4LLMs.
@@ 4.1.3 AI-intrinsic Threats and Defenses
AI intrinsic threats apply to LLMs due to the proximity of LLMs and AI, as shown in figure 2. Blockchain for AI (BC4AI) is an emerging technology, with blockchain-based solutions already being researched as a secure way to establish trust in the Internet of Things (IoT) [singh_convergence_2020][cuomo_how_2020]. Instead of the term BC4AI, some previous works refer to the integration as ”Onchain AI” [dinh_ai_2018][conway_opml_2024]. Research of BC4AI encapsulates other machine learning related techniques such as blockchain-based federated learning and blockchain for deep learning. Federated Learning is an addition to machine learning, as noted in figure 2, where federated learning uses a privacy-preserving and decentralized approach to centralized systems.
@@
There are multiple sources relating BC4AI, dating from 2018 to 2024  [dinh_ai_2018][wang_ai_2020][lopes_overview_2018] [witt_blockchain_2024][fan_blockchain_2024][salah_blockchain_2019][taherdoost_blockchain_2022][malhotra_blockchain-based_2024][calvaresi_explainable_2019][baranwal_somy_ownership_2019][bhumichai_convergence_2024]. Focusing on the most well known works, Salah et al. [salah_blockchain_2019] state integration benefits of BC4AI. For example, there are five main benefits, such as enhanced data security, improved trust in robotic decisions, collective decision-making, decentralized intelligence, and high efficiency[salah_blockchain_2019]. For enhanced data security, it is known that information held within the blockchain is highly secure. By storing sensitive and personal data in a distributed, disk-less environment and allowing AI algorithms to work to secure data, also to ensure more trusted and credible decision outcomes. The other benefits of improving trust within AI decision making involves using the blockchain as a record of the decision-making process, allowing better AI traceability to analyze the quality of responses. Secondly, from Dinh and Thai [dinh_ai_2018] who summarized the integration of blockchain and AI to where blockchain can assist AI in multiple aspects, as follows. AI can benefit in secure data sharing from blockchain, allowing transparency and accountability regarding which user’s data is accessed, when, and by whom, letting users maintain control of their personal data. Among other data concerns, with the integration of blockchain and AI, blockchain technologies can let user’s sell their data via smart contracts, enabling the possibility of data marketplaces without a centralized middleman, making the transactions private and secure between users.
A noteworthy framework from Malhotra et al. [malhotra_blockchain-based_2024] provides a blockchain-based proof-of-authenticity framework for explainable AI (XAI) utilizing a public Ethereum Blockchain, smart contracts, and IPFS (Interplanetary File System) to ensure secure, traceable, auditable transactions within the Ethereum network. This framework highlights three major components, smart contracts, an Ethereum and IPFS interconnected network, and a regulator, as depicted in figure 5. The use of smart contracts to enable continuous monitoring and tracing by all peers, in the case of any rule violations there are prompt rebound transactions to restore the system to an optimal state. To address the size limitation of storage on the blockchain, as further discussed in Section 6.1, Malhotra et al. [malhotra_blockchain-based_2024] used unique IPFS hashes stored on the Ethereum Blockchain to access larger-sized explanations that are stored off-chain in IPFS. These hashes are encrypted with the SHA256 algorithm to maintain data security. Thus, only entities with the corresponding hash can access and retrieve the IPFS hash and the associated explanation, ensuring controlled access even in a distributed network. Lastly, the regulator’s role is responsible for auditing and has access to the explanations to predict the user at fault using audit trails if system failure were to occur.

### v2

@@ 4 Existing Literature on BC4LLM
Independently, the fields of both LLMs and blockchain research have grown substantially over the past several years. It is no surprise that the literature surrounding these topics has begun to morph and relate to each other. In previous research, we have seen LLMs for Blockchain Security [he_large_2024] as well as an introduction to the term BC4LLM in Luo et al [luo_bc4llm_2023] where they provide a comprehensive survey of blockchain for LLMs. However, they do not acknowledge the multitude of safety and security solutions that blockchain provides for certain LLM vulnerabilities. Effectively, Luo et al. [luo_bc4llm_2023] aim to introduce BC4LLM for trusted AI, enabling reliable learning corpora, secure training processes, and identifiable generated content. In juxtaposition, our survey aims to analyze possible BC4LLM solutions closely related to our definitions of safety (2) and security (1) when looking at inherent system vulnerabilities in LLMs. To begin our analysis, we define these security problems based on previous work and highlight areas of research that are applicable to areas of BC4LLM safety and security.
4.1 Blockchain for LLMs’ Security
@@
Few papers and experiments analyze how the integration of these two technological powerhouses interacts with one another. We have seen benefits of this integration that apply to our definition of security (1). Balija et al. [balija_building_2024] introduce a peer-to-peer (P2P) federated LLM, namely PageRank, which works with a blockchain. This system operates in a fully decentralized capacity. Demonstrably, the blockchain implementation led to more efficient accuracy and latency results. With that being stated, Balija et al. [balija_building_2024] provide a developing direction in the field of BC4LLM to enhance system security. Below, we address several current vulnerabilities in LLMs and analyze them individually. In order to better understand these security problems, we categorize these vulnerabilities to their respective LLM training stages, highlight blockchain for AI works, and provide well-researched blockchain applications as a solution.
@@
Vulnerabilities are present at each step in the process of developing a LLM. In early methods of model training, we encounter adversarial attacks such as data poisoning and backdoor attacks within the corpus [shayegani_survey_2023, yan_protecting_2024]. Progressing into model fine-tuning and general use, the LLM can fall victim to prompt-based attacks [zhao_prompt_2023, abdali_securing_2024, yan_protecting_2024, liu_prompt_2024], inference attacks [keshk_privacy-preserving-framework-based_2020, hu_membership_2022], and RAG-related attacks [xue_badrag_2024, cheng_trojanrag_2024, anderson_is_2024, deng_pandora_2024, zou_poisonedrag_2024, zeng_good_2024]. These attacks are common vulnerabilities in both LLMs and AI since LLMs and AI are closely related as seen in Figure 2. Some of the threats against LLMs can be addressed by implementations from blockchain for AI (BC4AI) research, as elaborated below in Section 4.1.3. Considering the volume of potential attacks against LLMs, we make a further distinction of solutions that are specifically related to BC4LLM research and other blockchain-based solutions from BC4AI research. With this, we are able to highlight shared vulnerabilities for LLMs and AI. We provide an analysis of how blockchain can help defend against and mitigate these vulnerabilities, starting with threats during each phase of LLM training and utilization, continuing onto different blockchain solutions for AI inherent threats, and lastly noteworthy technology inherent attacks such as denial of service (DDoS) attacks and issues with supply chain logistics.
4.1.1 Blockchain for Threats in LLMs’ Training
@@
To mitigate the threats in LLMs’ training, data selection stands as a crucial aspect of model development, with particular emphasis on ensuring that training data is authentic, safe, and resistant to data poisoning attacks. One potential approach could be enabling the LLM to “unlearn” poisoned data or data deemed unsafe based on our definition of safety 2. Zuo et al. [zuo_federatedtrustchain_2024] establish federated TrustChain to enhance LLMs’ training and unlearning through a blockchain-based federated learning framework. Through integration with Hyperledger, the framework can efficiently perform unlearning, reducing the accuracy to 0.70% after unlearning given that the initial accuracy is 99.15% [zuo_federatedtrustchain_2024]. This demonstrates the potential of applying blockchain techniques to improve the security and privacy of LLMs, where LLMs can selectively forget specified data points while simultaneously preserving the performance via Low-Rank Adaptation (LoRA) and tuning hyper-parameters. This method of a blockchain-enabled federated unlearning process is further detailed as a future research possibility that has been thoroughly explored by few, as emphasized later in Section 7.1.
@@
Another significant issue is data poisoning. To address this issue, Gong et al. [gong_dynamic_2023] propose a possible blockchain solution, introducing dynamic large language models (DLLM) on blockchains. Instead of using the traditional centralized datasets that LLMs are provided with, developing LLMs on blockchains enables the creation of decentralized datasets. These datasets are less likely to be tampered with and can be easily audited for accuracy. Gong et al. [gong_dynamic_2023] present the DLLM to evolve after the training process. This was implemented by adjusting neural network parameters, enabling the LLM to continue learning during its use.
@@
Figure 4: An overview of a two-level framework, consisting of a privacy-based blockchain that uses the secure hash algorithm 512 (SHA512) to generate hashes for data integrity. Then, the enhanced proof of work (ePoW) is used to authenticate data records and prevent data poisoning attacks from altering original data. These hashes of data blocks are linked to each other, called a Hash Chain. Then, for the second level, a privacy-based variational autoencoder (VAE) for secure data transformation ensures robust protection against inference attacks while maintaining the utility model for anomaly detection.
@@
Additionally, blockchain-based systems can help assess where data poisoning may occur, and as shown in[keshk_privacy-preserving-framework-based_2020], blockchain can protect datasets and detect potential inference attacks through a two-level privacy preserving module. This research proposes a framework based on blockchain and deep learning, including two levels of privacy mechanisms as shown in Figure 4. For the first level, Keshk et al. [keshk_privacy-preserving-framework-based_2020] used SHA512 to generate secure hashes and then implemented an enhanced-proof-of-work (ePoW) technique for authenticating and preventing data poisoning attacks. The second level consisted of a VAE model for converting original data into an encoded format for mitigating inference attacks that could be learned from system-based machine learning. In their testing, these mechanisms were effective in preventing data poisoning and inference attacks from manipulated smart power network datasets. BC4LLMs could benefit from this similar type of implementation, working with secure methods of hashing and blockchain-based deep learning privacy preservation techniques. By integrating a two-level privacy-preserving module, BC4LLMs can ensure data integrity and confidentiality while effectively detecting and mitigating both data poisoning and inference attacks.
@@
Poisoned data has been an interest with RAG in particular. For example, Xue et al. [xue_badrag_2024] develop a way to identify security vulnerabilities from a poisoned corpus, but they do not use blockchain as a solution. We address the absence of research on blockchain and RAG, especially when using blockchain to help prevent RAG security issues. We discuss this as a possible future research direction as there remains a current gap in research of blockchain-based RAG systems and elaborate on this topic in Section 7.2. Poisoned data overall is a major concern within LLMs and we offer blockchain as a potential source of ground truth to aid in mitigating this threat during the pre-training stages of LLMs and potentially mitigate RAG security concerns. In addition to data poisoning, LLMs are susceptible to backdoor attacks hidden in the training data during the LLM pre-training phase. Zhao et al. [zhao_prompt_2023] introduce ProAttack which improves the stealth of backdoor attacks by accurately labeling poisoned data samples. As these attacks improve and become more sophisticated, it is crucial to explore robust defense mechanisms for LLMs. Few defense mechanisms using blockchain techniques have been studied, while Li et al. [li_blockchain-based_2023] propose a blockchain-based federated-learning framework (DBFL) that withstands backdoor attacks in a blockchain environment by incorporating an RLR aggregation strategy into the aggregation algorithm of a user and the addition of gradient noise to limit the effectiveness of backdoor attacks. The robustness of FL against backdoor attacks is enhanced by using various blockchain functions, including digital signature verification and simulation of chain resynchronization [li_blockchain-based_2023].
4.1.2 Blockchain for Threats in LLMs’ Prompting and Utilization
@@
LLMs are often further trained through techniques such as instruction tuning, alignment tuning, and fine-tuning, each of which may introduce specific vulnerabilities, such as prompt injection [mbula_mboma_assessing_2023, shen_large_2023] and backdoor attacks [li_blockchain-based_2023]. These adversarial attacks are included under the term active privacy 3 where a malicious user attempts to gain unauthorized access. Blockchain technology, with its inherent transparency and immutability, holds the potential to mitigate and defend against these vulnerabilities. For instance, blockchain’s ability to ensure data integrity and provide traceability can play a pivotal role in defending against prompt injection attacks. Mbula et al. [mbula_mboma_assessing_2023] produce an overview of LLMs for blockchain and note the capabilities of blockchain, especially the transparency and immutability to provide a reliable audit trail of transactions to track and investigate any suspicious activities. Although not specifically focused on prompt injection, this approach showcases how blockchain can supplement security by offering a clear and immutable record of interactions. Applying this to BC4LLMs can help prevent suspicious users from continuously interacting with an LLM, allowing for traceability to stop the user from entering malicious prompts. We recognize prompt injection is a critical vulnerability in LLMs and AI-related systems, yet as noted in the survey from Shen et al. [shen_large_2023] few blockchain defenses for prompt injection are present in the current field of research.
@@
Inference attacks are also a critical concern for LLMs and active privacy, as malicious users may attempt to extract sensitive data from the model, hence why the Taxonomy 3 has inference attacks standalone. As discussed previously in Section 4.1.1, Keshk et al. [keshk_privacy-preserving-framework-based_2020] apply Blockchain and DL techniques to preserve privacy and prevent inference attacks through a framework as depicted in Figure 4. For more inference attack applicable work, a survey [hu_membership_2022] thoroughly discusses membership inference attacks on ML and provides a group of defenses including differential privacy, regularization, confidence masking, and knowledge distillation. In other related works, there are instances of blockchain-based differential privacy methods [zhao_blockchain-based_2021, han_blockchain-based_2020, park_blockchain-based_2021], but current research that uses blockchain-based differential privacy frameworks to prevent inference attacks is limited; It is worth noting that the theoretical foundation and potential synergies of this combination are promising. Another area with limited research is blockchain as a defense for jailbreaking attacks, which exploit the inherent capabilities of LLMs to bypass restrictions. There are multiple articles defending LLMs from jailbreaking attacks, yet little to none fully include blockchain to prevent jailbreaking. Hu et al. [hu_blockchain-based_2021] explores a blockchain defense mechanism for malware checking on operating systems, indicating a possible direction for future research in integrating blockchain to defend against jailbreaking in LLMs. As previously explained in Section 3.2, backdoor attacks after the model has been trained are based on modifications to the trained model. The key blockchain-based federated-learning framework from Li et al. [li_blockchain-based_2023] discussed in detail in Section 4.1.1 used a combination of a blockchain environment and an RLR aggregation strategy to defend against backdoor attacks. This framework effectively coordinated FL processes and maintained learning security and user privacy. When testing backdoor attacks caused by malicious participants, the accuracy of the model increased when using the RLR aggregation strategy [li_blockchain-based_2023]. Given these findings, the possibility of leveraging blockchain transparency and immutability presents a robust mechanism for improving LLM security against active privacy threats. However, comprehensive integration and empirical validation of blockchain-based defenses in LLMs remain imperative to advance the field of BC4LLMs.
@@ 4.1.3 AI-intrinsic Threats and Defenses
AI intrinsic threats apply to LLMs due to the proximity of LLMs and AI, as shown in Figure 2. Blockchain for AI (BC4AI) is an emerging technology, with blockchain-based solutions already being researched as a secure way to establish trust in the Internet of Things (IoT) [singh_convergence_2020, cuomo_how_2020]. Before BC4AI, some previous works refer to the integration as “Onchain AI” [dinh_ai_2018, conway_opml_2024]. Research of BC4AI encapsulates other machine learning techniques, such as blockchain-based federated learning and blockchain for deep learning. Federated Learning is an addition to machine learning, as noted in Figure 2, where federated learning uses a privacy-preserving and decentralized approach to centralized systems.
@@
A substantial literature on BC4AI has emerged from 2018 to 2024 [dinh_ai_2018, wang_ai_2020, lopes_overview_2018, witt_blockchain_2024, fan_blockchain_2024, salah_blockchain_2019, taherdoost_blockchain_2022, malhotra_blockchain-based_2024, calvaresi_explainable_2019, baranwal_somy_ownership_2019, bhumichai_convergence_2024]. Among the most notable works, Salah et al. [salah_blockchain_2019] state the integration benefits of BC4AI. For example, there are five main benefits, such as enhanced data security, improved trust in robotic decisions, collective decision-making, decentralized intelligence, and high efficiency [salah_blockchain_2019]. For enhanced data security, information stored within a blockchain is considered highly secure. By storing sensitive and personal data in a distributed, disk-less environment, blockchain can work alongside AI algorithms to strengthen data protection and promote more trusted and credible decision outcomes. The other benefits of improving trust within AI decision-making involve using the blockchain as a record of the decision-making process, allowing better AI traceability to analyze the quality of responses. Secondly, Dinh and Thai [dinh_ai_2018] summarize the integration of blockchain and AI to where blockchain can assist AI in multiple aspects, as follows. AI can benefit in secure data sharing from blockchain, allowing transparency and accountability regarding which user’s data is accessed, when, and by whom, letting users maintain control of their personal data. Among other data concerns, with the integration of blockchain and AI, blockchain technologies allow users to trade data via smart contracts, enabling the possibility of data marketplaces without a centralized middleman, making the transactions private and secure between users. Besides, Malhotra et al. [malhotra_blockchain-based_2024] propose a blockchain-based proof-of-authenticity framework for explainable AI (XAI) utilizing a public Ethereum Blockchain, smart contracts, and IPFS (Interplanetary File System) to ensure secure, traceable, auditable transactions within the Ethereum network. This framework highlights three major components, smart contracts, an Ethereum and IPFS interconnected network, and a regulator, as depicted in Figure 5. Using smart contracts can enable continuous monitoring and tracing by all peers, in the case of any rule violations there are prompt rebound transactions to restore the system to an optimal state. To address the size limitation of storage on the blockchain, as further discussed in Section 6.1, Malhotra et al. [malhotra_blockchain-based_2024] apply unique IPFS hashes stored on the Ethereum Blockchain to access larger-sized explanations that are stored off-chain in IPFS. These hashes are encrypted with the SHA256 algorithm to maintain data security. Thus, only entities with the corresponding hash can access and retrieve the IPFS hash and the associated explanation, ensuring controlled access even in a distributed network. Lastly, the regulator’s role is responsible for auditing and has access to the explanations to predict the user at fault using audit trails if system failure were to occur.